mom summary 19 01 10 n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
MOM Summary 19/01/10 PowerPoint Presentation
Download Presentation
MOM Summary 19/01/10

Loading in 2 Seconds...

play fullscreen
1 / 9

MOM Summary 19/01/10 - PowerPoint PPT Presentation


  • 99 Views
  • Uploaded on

Water Now the ice has cleared I am able to get on the roof again. Could do with getting the conductivity readings fed back similarly. I believe A.Nichols has approached J. Governs wrt this – I haven’t heard anything yet.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'MOM Summary 19/01/10' - jacoba


Download Now An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
mom summary 19 01 10

Water

    • Now the ice has cleared I am able to get on the roof again.
    • Could do with getting the conductivity readings fed back similarly. I believe A.Nichols has approached J. Governs wrtthis – I haven’t heard anything yet.
    • The water level appears to keep dropping – I’m assuming that it is being topped up by J.Governs. No evidence of any leaks though?
    • Magnet Chillercontroller outside of the MLCR has been fixed to prevent it crashing on a daily basis

MOM Summary 19/01/10

mom summary

MICE HALL

  • Decay Solenoid
    • The New collar was fitted last week, Transfer line vacuum appears to have improved as a result.
    • Scaffold was put up for Linde last week.
    • Servicing due this week.
    • It will be necessary to start cooldown immediately to ensure DS is ready for machine physics.
  • CryoCompressers
    • Unsure of status of these.
  • Other Work
    • Willie’s team are continuing to drill holes in the steelwork in preparation for the tracker solenoids.

MOM Summary

mom summary1

Daresbury

    • Guys came down on the 12th 13th Jan.
    • Helped to installLuminosity Monitors in target Synch.
    • Brought replacementcomponent for target PSU
    • Installednew interlock for target frame (will not be functional yet though) – but there is an issue with a return signal to be solved.
    • There is still some issue with TC cabling numbers to the target to be resolved.
    • (Fixed) an issue with the beam-on light on the upper DS Gate (I think it is now permanently off!)
    • IM returned on Mon 18th
    • Look at return signal on target interlock.
    • Wiring of power to TOF 2.
    • Take a closer look at TC cabling to the target.
    • Anything ELSE?

MOM Summary

mom summary2

Luminosity Monitors

    • Paul Soler came down Tue to Fri last week to install LM
    • Cabling pulled through by Daresbury guys
    • Electronics is in the bottom third of rack next to target.
    • One PMT has failed, Paul may be able to get a replacement for machine physics but it doesn’t compromise operation. My understanding is that it may just increase the S/N ratio slightly.
  • DAQ
    • As far as I am aware there is no change - He plans to come 26th to 31st of Jan to fix CAMAC. I have requested documentation so that a non-daq expert can plug leads back in.

MOM Summary

mom summary3

Lab 7

    • Helium leak developed in Lab 7, source not yet found. This emptied the cylinder
    • Due to planned power cut in R1, Lab 7 was effectively switched off on Friday 15th. I believe the plan is to get everything working again in preparation for the trackers moving to the MICE hall.
  • Power Cut
    • The power was cut in the MICE hall on Saturday morning (This was a planned cut to allow the power to be switched to a new distribution board)
    • Unfortunately this also affected the MLCR (not forseen) and several computers were affected
    • The DS vacuums had been previously isolated but without the computers the vacuum pumps on the turret could not be restrarted until Monday.

MOM Summary

mom summary4

Running

    • No running planned for January because of shutdown.
    • Machine Physics to run from 2nd to 14th February.
    • We have informally requested 5 shifts, formal confirmation and shift dates tbc
    • I have NO plan for User run. What would we like to achieve during this period?
  • Rehersal
    • There will be a rehersal on the afternoon of the 28th of Jan to ensure that all equipment appears to be functional. Simultaneously Beam Line training will be given to beam line experts
    • This will be preceded by a meeting in the morning (10:30) to determine the ‘Beam Line Expert’ rota

MOM Summary

mom summary5

Shift 1:   Friday 5th February 16:30 - 00:00Shift 2: Sat 6th Feb OR Sun 7th Feb. This has yet TBC by Dean.  This could be a 16h shift if we can utilise the time. If we can, then we have got a shift for free! Thanks DeanShift 3: Mon 8th February 16:30 - 00:00Shift 4 & 5: Wed 10th February. This is for the 16h activation run at 2V beamloss.  09:00 - FinishThere may be the possibility of additional shifts after the 10th but this is TBC.

MOM Summary

mom summary6

NOTE : All TBC!

  • Shift 1: Objectives:
    • Re-establish DAQ and target operations
    • Reset and setup all Detectors (GVA1 / TOF1 & 2/ KL)
    • Re-establish Online Reconstruction
    • Commission Luminosity Monitors (Parasitic)
  • Shift 2 & 3
    • Commission Luminosity Monitors
    • Calibrate FNAL1 monitor and GVA1 monitor
    • Adams Beamlossvs rate studies

MOM Summary

mom summary7

Shift 4 & 5 (16h Combined)

    • Establish Target Operation at Required Beamloss of 2V
    • Ensure all Detectors are Running
    • Run for 16 Hours
    • If target insertion rate is 1 every 2.56 seconds then the total number of actuations in 16h is 22,500. Therefore there will need to be at least 4 calibration runs! (One every 3.5 hours)

MOM Summary